Calculate the pH of the following solutions: 

0.05M HCl

0.65M NaOH


1
Expert's answer
2021-03-29T06:08:29-0400

Since it is a strong acid and it fully dissolves , the concentration of HCl equals to the [H+] (H3O+) concentration.

Hence, (H+)= 0.05M

pH-log[H+]

pH=-log(0.05)

PH of 0.05M HCl =1.3

NaOH is a strong base, meaning it will dissociate basically completely in solution to form Na+ and OH− ions in a solution.

POH = - log (0.65)

= 0.2

PH + POH = 14

= 14-0.2

PH of 0.65 M NaOH = 13.8